OpenStudy (wwe123):

If the sum of first n terms of two arithmetic progression are in the ratio (7n -5 ) : (5n + 17), show that the 6th terms of the two series are equal

OpenStudy (wwe123):

??

OpenStudy (asnaseer):

hint: find out the value of n where the ratio of the sums of the two series are equal

OpenStudy (asnaseer):

@wwe123 do you understand?

OpenStudy (cwrw238):

ah yes !! i must admit i missed that..

OpenStudy (asnaseer):

:) it is /subtle/

OpenStudy (wwe123):

this i known

OpenStudy (asnaseer):

so what value of n do you get for this?

OpenStudy (wwe123):

i am not able to solve the equation

OpenStudy (asnaseer):

find the value of n that satisfies: (7n -5 ) = (5n + 17)

OpenStudy (wwe123):

how (7n -5 ) = (5n + 17)

OpenStudy (asnaseer):

that would make the ratio of the sums of both series equal

OpenStudy (asnaseer):

once you have found n, then use the formula for the sum of the first n terms of a series to get a ratio of the two series at that value of n. you will then see the desired answer to the question "pop" out

OpenStudy (wwe123):

how

OpenStudy (asnaseer):

try it yourself and you will see

OpenStudy (asnaseer):

if you list your steps here then I can guide you further

OpenStudy (wwe123):

one min

OpenStudy (wwe123):

\[\frac{ 2a _{1}+(n-1)d _{1} }{ 2a _{2}+(n-1)d _{2} }=\frac{ 7n-5 }{ 5n-17 }\]

OpenStudy (wwe123):

this is right ^^^^

OpenStudy (asnaseer):

yes - but first find the value of n that makes this ratio equal to 1

OpenStudy (wwe123):

how to find the value of n ?

OpenStudy (asnaseer):

look at the right-hand-side of your equation - set it to 1 and solve for n

OpenStudy (asnaseer):

i.e. solve this:\[\frac{ 7n-5 }{ 5n-17 }=1\]

OpenStudy (wwe123):

but how RHS =1

OpenStudy (asnaseer):

I am postulating that if we find a value for n that makes this ratio equal to 1, then we will be able to get somewhere towards the answer.

OpenStudy (wwe123):

ok

OpenStudy (wwe123):

i amnot able to solve, plz show

OpenStudy (asnaseer):

surely you can solve this:\[\frac{ 7n-5 }{ 5n-17 }=1\]

OpenStudy (asnaseer):

multiply both sides by 5n-17

OpenStudy (asnaseer):

and solve for n this is basic algebra

OpenStudy (wwe123):

n= 11

OpenStudy (asnaseer):

good - now use this value of n in your original equation:\[\frac{ 2a _{1}+(n-1)d _{1} }{ 2a _{2}+(n-1)d _{2} }=\frac{ 7n-5 }{ 5n-17 }\]

OpenStudy (asnaseer):

to simplify it - what do you end up with?

OpenStudy (wwe123):

there are 4 unknown value how cani solve \[a _{1},a _{2},d _{1}, d _{2}\]

OpenStudy (asnaseer):

do not solve - just simplify the expression by substituting n=11 into it

OpenStudy (wwe123):

\[\frac{ 2a _{1}+10d _{1} }{ 2a _{2}+10d _{2} }=\frac{ 72 }{ 38 }\]

OpenStudy (wwe123):

@asnaseer

OpenStudy (asnaseer):

the RHS is incorrect

OpenStudy (asnaseer):

and you can divide the LHS by 2 to simplify the fraction further

OpenStudy (asnaseer):

I see the mistake - it should be 5n+17

OpenStudy (asnaseer):

not 5n-17

OpenStudy (asnaseer):

\[\frac{ 2a _{1}+(n-1)d _{1} }{ 2a _{2}+(n-1)d _{2} }=\frac{ 7n-5 }{ 5n+17 }\]

OpenStudy (wwe123):

so ,rhs=0

OpenStudy (asnaseer):

no???

OpenStudy (asnaseer):

if n=11 what is 7n-5=?

OpenStudy (wwe123):

sorry its 1

OpenStudy (asnaseer):

correct, so you have ended up with:\[\frac{ 2a _{1}+10d _{1} }{ 2a _{2}+10d _{2} }=1\]now simplify the fraction on the LHS (you can divide numerator and denominator by 2)

OpenStudy (asnaseer):

and then notice what each term (i.e. numerator and denominator) represent

OpenStudy (wwe123):

i done it ...................... :)

OpenStudy (asnaseer):

good :)

OpenStudy (asnaseer):

so you see how the answer just "popped out" :)

OpenStudy (wwe123):

yeah